OpenStudy (anonymous):

3. Having survived the meteor impact, thanks to some last minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y^2 + x^2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’s path.

OpenStudy (anonymous):

@woohoo

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

i need serious help!

OpenStudy (phi):

y^2 + x^2 = 40,000 is the equation of a circle, centered at (0,0) , with radius = 200 (notice 200*200= 40000) they say Using the original trajectory of the ship I assume you have an equation ? we need it to answer this question

OpenStudy (anonymous):

my equation is y=5x+1 @phi

OpenStudy (phi):

the problem is asking where your line intersects the circle.

OpenStudy (anonymous):

Im still lost i don`t get it at all @phi

OpenStudy (phi):

Here is a graph of the problem

OpenStudy (phi):

you have two equations y=5x+1 y^2 + x^2 = 40,000 notice that if you replace y in the second equation using y= 5x+1 you get \[ (5x+1)^2 +x^2 = 40000 \] first expand (5x+1)(5x+1) can you do that ?

OpenStudy (anonymous):

which means i have to use foil? @phi

OpenStudy (phi):

yes

OpenStudy (anonymous):

so i got 25^2 + 10x + 5x + 1 ?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

F(irst) 5x*5x or 25x^2 O(uter) 5x*1 or 5x I(nner) 1*5x or 5x L(ast) 1*1 or 1 if you combine the 5x+5x to 10x we get 25x^2 +10x+1 so you are close

OpenStudy (phi):

you now have \[ (5x+1)^2 +x^2 = 40000 \\ (5x+1)(5x+1) +x^2 = 40000 \\ 25x^2 +10x+1 + x^2 = 40000 \] now combine the "like terms". In this case the x^2 terms. also add -40000 to both sides. can you do that ?

OpenStudy (anonymous):

so with the like terms with x^2 and 25x^2 which give you 26x^2 @phi

OpenStudy (phi):

yes. You can think: 25 x squarers plus one more x squared is 26 x squarers (I think of x^2 as a thing we count, and we have 26 of them) \[ 25x^2 +10x+1 + x^2 = 40000 \\26x^2 +10x+1 = 40000 \] now add -40000 to both sides.

OpenStudy (phi):

to add -40000 to both sides you write -40000 on both sides

OpenStudy (anonymous):

both sides? @phi

OpenStudy (anonymous):

im not sure if i did it right because i subtract 40000 from 10x and got 1,5.38

OpenStudy (phi):

yes. If you have an equation, it will stay an equation if you add the same thing to both sides. (to keep things fair, so to speak)

OpenStudy (phi):

when you add letters and numbers, you don't get *just* numbers. Here, to add -40000 to both sides we do this \[ 26x^2 +10x+1-40000 = 40000 -40000 \]

OpenStudy (phi):

notice we added -40000 to both sides. now simplify. and you can only combine *like terms* so the x terms don't get involved.

OpenStudy (anonymous):

im confused @phi

OpenStudy (phi):

\[ 26x^2 +10x+1-40000 = 40000 -40000 \] what is 1-40000 ? on the left side ?

OpenStudy (anonymous):

39,999 ? @phi

OpenStudy (phi):

the correct numbers, but the answer is negative: -39999

OpenStudy (phi):

what about the right hand side of the = sign ? what is 40000 - 40000 ?

OpenStudy (anonymous):

0? @phi

OpenStudy (phi):

yes. so now you have \[ 26x^2+10x-39999=0 \] that is a "quadratic equation" (because it has one variable , the x, and the highest exponent on the x is a 2 to solve for x, you use the quadratic formula. It is more complicated than the algebra we have done so far, so you probably should watch Khan's video (maybe a few times!) and then try using it. You will need a calculator. You should get the same numbers as in the graph I posted up above. http://www.khanacademy.org/math/algebra/quadratics/quadratic_formula/v/using-the-quadratic-formula